OpenStudy (anonymous):

4 7/12 + 2 3/4 + 7 5/12

OpenStudy (anonymous):

First of all, this will be easiest if all the denominators are the same, so let's change 4 to 12. making the problem, 4 7/12 + 2 9/12 + 7 5/12.

OpenStudy (ageta):

Answer: = 15/2 + 89/12 = ((15 × 12) + (89 × 2)) / (2 × 12) = (180 + 178) / 24 = 358/24 = 179/12 = 14 11/12

OpenStudy (anonymous):

From here you will probably want to make them an improper fraction,

OpenStudy (12man):

\[14\frac{ 11 }{ 12 }\]

OpenStudy (anonymous):

55/12 + 35/12 + 89/12=

OpenStudy (ageta):

yes it it 14 11/12

OpenStudy (lizz123):